OpenStudy (anonymous):

5. If f(x) = 2x − 2 and g(x) = 3x − 1, find (f − g)(4).

OpenStudy (anonymous):

Do the same thing, but this time, subtract f(x) and g(x)

OpenStudy (anonymous):

so can you set it up for me...thats the problem with these equations

OpenStudy (anonymous):

Whats F(x)?

OpenStudy (anonymous):

2x − 2

OpenStudy (anonymous):

Now subtract g(x)

OpenStudy (anonymous):

so 2x

OpenStudy (anonymous):

@marvin.cling

OpenStudy (anonymous):

Nope, try it again. oure not combining ht elike terms correctly

OpenStudy (anonymous):

So 4x?

OpenStudy (anonymous):

No.. Do this: 2x − 2-( 3x − 1) Distribute the negative sign: 2x-2-3x+1 Now do the combining like terms: 2x-3x and -2+1

OpenStudy (anonymous):

\[f-g=(2x-2)-(3x-1)=-x-1,(4)=-5\]

OpenStudy (anonymous):

i gotta sign off now gl with u hw

OpenStudy (anonymous):

@marvin.cling Youre not supposed to give the answer, you're supposed to help, step by step